### v3

# Variational inequalities (VIs) have attracted researchers’ attention in various fields for more than half a century [Browder, 1965]. In this work, we investigate the variational inequality problem in the following form:

BRAIn Lab[Uncaptioned image] Shuffling Heuristic in Variational Inequalities: Establishing New Convergence Guarantees Daniil Medyakov1,2111Corresponding author: medyakovd3@gmail.com, Gleb Molodtsov1,2, Grigoriy Evseev1, Egor Petrov1, Aleksandr Beznosikov1,2 1Basic Research of Artificial Intelligence Laboratory (BRAIn Lab) 2Federated Learning Problems Laboratory Variational inequalities have gained significant attention in machine learning and optimization research. While stochastic methods for solving these problems typically assume independent data sampling, we investigate an alternative approach – the shuffling heuristic. This strategy involves permuting the dataset before sequential processing, ensuring equal consideration of all data points. Despite its practical utility, theoretical guarantees for shuffling in variational inequalities remain unexplored. We address this gap by providing the first theoretical convergence estimates for shuffling methods in this context. Our analysis establishes rigorous bounds and convergence rates, extending the theoretical framework for this important class of algorithms. We validate our findings through extensive experiments on diverse benchmark variational inequality problems, demonstrating faster convergence of shuffling methods compared to independent sampling approaches.
1  Introduction
Variational inequalities (VIs) have attracted researchers’ attention in various fields for more than half a century [Browder, 1965]. In this work, we investigate the variational inequality problem in the following form:
$\text{find}\penalty 10000\ \penalty 10000\ z^{*}\in Z\penalty 10000\ \penalty 10000\ \text{such that}\penalty 10000\ \penalty 10000\ \forall z\in Z\hookrightarrow\langle F(z^{*}),z-z^{*}\rangle+g(z)-g(z^{*})\geqslant 0,$ (1)
where $F$ is a monotone operator and $g$ is a proper convex lower semicontinuous function, which serves as a regularizer. Variational inequalities serve as a universal tool for addressing specific problems, such as minimization, saddle point problems, fixed point problems, and others [Facchinei and Pang, 2003; Kinderlehrer and Stampacchia, 2000]. We provide examples to offer intuition about VIs.
Example 1 (Convex optimization).
Consider the following convex regularized optimization problem:
$\underset{z\in\mathbb{R}^{d}}{\min}\left[f(z)+g(z)\right].$ (2)
In this example, $f$ is a smooth data representative term, and $g$ is probably a non-smooth regularizer. We define $F(z)=\nabla f(z)$. Then $z^{*}\in\text{dom}\penalty 10000\ g$ is the solution of (1) if and only if $z^{*}\in\text{dom\penalty 10000\ }g$ is the solution of (2). In this way, the problem (2) can be considered as a variational inequality.
Example 2 (Convex-concave saddles).
Consider the following convex-concave saddle point problem:
$\underset{x\in\mathbb{R}^{d_{x}}}{\min}\underset{y\in\mathbb{R}^{d_{y}}}{\max}\left[f(x,y)+g_{1}(x)-g_{2}(y)\right].$ (3)
There, $f$ has the same interpretation as in Example 1, and $g_{1}$, $g_{2}$ can also be perceived as regularizers. We define $F(z)=F(x,y)=\left[\nabla_{x}f(x,y),-\nabla_{y}f(x,y)\right]$. Then $z^{*}\in\text{dom}\penalty 10000\ g_{1}\times\text{dom}\penalty 10000\ g_{2}$ is the solution of (1) if and only if $z^{*}\in\text{dom}\penalty 10000\ g_{1}\times\text{dom}\penalty 10000\ g_{2}$ is the solution of (3). In this way, the problem (3) can be considered as a variational inequality.
There are multiple practical reasons to focus on this formulation. Firstly, for numerous non-smooth problems, solutions are obtained more efficiently when formulated as saddle point problems [Nesterov, 2005; Nemirovski, 2004; Chambolle and Pock, 2011; Esser et al., 2010]. Secondly, recent studies have established new connections between VIs and reinforcement learning [Omidshafiei et al., 2017; Jin and Sidford, 2020], adversarial training [Madry et al., 2017], and generative adversarial networks (GANs) [Goodfellow et al., 2014]. In particular, consideration of monotone and strongly monotone inequalities provides useful methods and recommendations for the GAN community [Daskalakis et al., 2017; Gidel et al., 2018; Mertikopoulos et al., 2018; Chavdarova et al., 2019; Liang and Stokes, 2019; Peng et al., 2020]. VIs also have extensive applications in various classic problems, including discriminative clustering [Xu et al., 2004], matrix factorization [Bach et al., 2008], image denoising [Esser et al., 2010; Chambolle and Pock, 2011], robust optimization [Ben-Tal et al., 2009], economics, game theory [Von Neumann and Morgenstern, 1953], and optimal control [Facchinei and Pang, 2003].
Solving the problem (1) requires specialized methods, as traditional optimization techniques, e.g. the Gradient Descent method, often fall short when applied to VIs and saddle point problems [Harker and Pang, 1990]. These classic methods not only struggle with efficiency but also provide weak theoretical convergence guarantees in the VI context [Beznosikov et al., 2023]. Among the various approaches developed for VIs, the Extragradient method [Korpelevich, 1976; Mokhtari et al., 2020] stands out as one of the most fundamental and effective techniques.
While variational inequalities provide a powerful framework for addressing a wide range of problems, recent trends in machine learning and data science present new challenges. The exponential growth in dataset sizes and the increasing complexity of models create a pressing need for more efficient computational approaches [Bottou, 2010; Dean et al., 2012; Medyakov et al., 2023]. To address these challenges in the context of VIs, we reformulate the problem by considering the operator $F$ as the finite sum of operators $F_{i}$:
$F(z)=\frac{1}{n}\sum\nolimits_{i=1}^{n}F_{i}(z),$ (4)
where each $F_{i}$ corresponds to an individual data point. This decomposition enables us to address large-scale problems more effectively.
In this paper, we explore stochastic algorithms that are particularly suitable for practical extensive applications. As mentioned previously, the number of operators $n$ is typically large, making the computation of the full operator value at each iteration computationally expensive. Instead, stochastic algorithms randomly select $F_{i}$ at each iteration. The stochastic version of the Extragradient method [Juditsky et al., 2011] selects random independent indices $i_{t},j_{t}$ at iteration $t$ and performs the following updates:
$z^{t+\frac{1}{2}}$ $=z^{t}-\gamma F_{i_{t}}(z^{t}),$
$z^{t+1}$ $=z^{t}-\gamma F_{j_{t}}(z^{t+\frac{1}{2}}).$
Just as deterministic Extragradient is a modification of the classic gradient method with an additional step, the stochastic Extragradient represents the same modification of SGD [Robbins and Monro, 1951]. Although this method performs well on variational inequalities, it encounters a significant issue with its properties and performance that has been thoroughly studied: the variance of its inherent stochastic estimators of operators remains high throughout the learning process. Hence, Extragradient with a constant learning rate converges linearly only to a neighborhood of the optimal solution, the size of which is proportional to the step size and variance [Juditsky et al., 2011]. This problem is also characteristic of classic SGD [Bottou, 2009; Moulines and Bach, 2011; Gower et al., 2020].
To address this limitation, the variance reduction (VR) technique has been developed for a classic finite-sum minimization task [Johnson and Zhang, 2013]. The method involves the following steps: at the $t$-th iteration, an index $i_{t}$ is selected along with a reference point $\omega^{t}$, which is updated once per epoch or selected probabilistically (as in loopless versions, e.g., [Kovalev et al., 2020]). Considering the convex optimization problem (see Example 1), we can formally write the stochastic reduced gradient at the point $z^{t+\frac{1}{2}}$ as
$\nabla\hat{f}_{i_{t}}(z^{t+\frac{1}{2}})=\nabla f_{i_{t}}(z^{t+\frac{1}{2}})-\nabla f_{i_{t}}(\omega^{t})+\nabla f(\omega^{t}).$
The objective of the variance reduction mechanisms is to overcome the limitations of naive gradient estimators. These mechanisms employ an iterative process to construct and apply a gradient estimator with progressively reduced variance. This approach permits the use of larger learning rates, thereby accelerating the training process.
Besides, along with the aforementioned SVRG, there are popular methods for solving the classic finite-sum problem based on this technique, such as SAG [Roux et al., 2012], SAGA [Defazio et al., 2014a], Finito [Defazio et al., 2014b], SARAH [Nguyen et al., 2017; Hu et al., 2019], and SPIDER [Fang et al., 2018]. The variance reduction mechanism is utilized not only in methods that address the minimization problem but also in methods for the problem (1). Examples include the VR versions of Extragradient, Mirror-prox [Alacaoglu and Malitsky, 2022], gradient method [Palaniappan and Bach, 2016], and forward-reflected-backward (FoRB) [Alacaoglu et al., 2021].
In addition to stochastic methods, various heuristics for selecting the $i_{t}$-th index at each iteration of algorithms exist. Careful analysis of these strategies may lead to the development of more robust and efficient algorithms. In this paper, we explore the shuffling heuristic [Mishchenko et al., 2020a; Safran and Shamir, 2020; Koloskova et al., 2023; Malinovsky et al., 2023]. Unlike the random and independent selection of the index $i_{t}$ at each iteration, which is common in classic stochastic methods, this heuristic adopts a more practical approach. Specifically, it involves permuting the sequence of indices $\{1,\ldots,n\}$, where $n$ is the number of data samples (4), and then selecting the index corresponding to the iteration number during the algorithm’s execution. This approach ensures that during one epoch of training, we take a step for each operator, and only once. There are several shuffling techniques available. Among the most popular are Random Reshuffling (RR) [Gürbüzbalaban et al., 2021; Haochen and Sra, 2019; Nagaraj et al., 2019], where data is shuffled before each epoch; Shuffle Once (SO) [Safran and Shamir, 2020; Rajput et al., 2020], where shuffling occurs once before the start of training; and Cyclic permutation [Mangasarian and Solodov, 1993; Bertsekas and Tsitsiklis, 2000; Nedic and Bertsekas, 2001; Li et al., 2019], where data is accessed deterministically in a cyclic order.
Related Works. There are many methods available to solve the problem of variational inequalities. As mentioned above, the standard deterministic choice for solving the problem (1) is Extragradient [Korpelevich, 1976]. This method addresses variational inequalities in the Euclidean setup. Later, Mirror-prox [Nemirovski, 2004], which exploits the Bregman divergence, was proposed. This approach accounts for generalized geometry that may be non-Euclidean. Additionally, there is a set of deterministic methods for solving variational inequalities: forward-backward-forward (FBF) [Tseng, 2000], Dual extrapolation [Nesterov, 2007], reflected gradient [Malitsky, 2015], forward-reflected-backward (FoRB) [Malitsky and Tam, 2020].
For the first time, the stochastic version of algorithms for solving VIs was proposed in the work [Juditsky et al., 2011]. Later, to reduce the variance inherent in these stochastic methods, researchers adopted variance reduction techniques. The initial works in this field are [Palaniappan and Bach, 2016; Chavdarova et al., 2019]. In particular, [Palaniappan and Bach, 2016] studied the stochastic gradient method with variance reduction. The method was based on SVRG [Johnson and Zhang, 2013] and incorporated Catalyst envelope acceleration. The combination of Extragradient and SVRG was considered in [Chavdarova et al., 2019]. They achieved a worse convergence rate compared to [Palaniappan and Bach, 2016] and only in the strongly monotone setting. Consequently, a notable paper in which the authors considered monotone operators was presented [Carmon et al., 2019]. This work also falls under the Bregman setup but requires additional assumptions on the operator $F$ and considers the matrix games setup. The current state-of-the-art in this area is the article [Alacaoglu and Malitsky, 2022], which improved the convergence estimates of all previous studies. This work addressed various scenarios, including generally monotone and strongly monotone operators, as well as the Bregman and Euclidean setups. Convergence results from the papers highlighted above are summarized in Table 1.
Algorithm Sampling VR?
Strongly
Monotone
Complexity
Monotone
Complexity
Extragradient [Korpelevich, 1976; Mokhtari et al., 2020] Deterministic $\mathcal{\widetilde{O}}\left(\frac{nL}{\mu}\right)$ $\mathcal{O}\left(\frac{nL}{\varepsilon}\right)$
Mirror-prox [Nemirovski, 2004] Deterministic $\backslash$ $\mathcal{O}\left(\frac{nL}{\varepsilon}\right)$
FBF [Tseng, 2000] Deterministic $\backslash$ $\mathcal{O}\left(\frac{nL}{\varepsilon}\right)$
FoRB [Malitsky and Tam, 2020] Deterministic $\backslash$ $\mathcal{O}\left(\frac{nL}{\varepsilon}\right)$
Mirror-prox [Juditsky et al., 2011] Independent $\backslash$ $\mathcal{O}\left(\frac{L}{\varepsilon}+\frac{1}{\varepsilon^{2}}\right)$
Extragradient [Beznosikov et al., 2020] Independent $\mathcal{\widetilde{O}}\left(\frac{L}{\mu}+\frac{1}{\mu^{2}\varepsilon}\right)$ $\mathcal{O}\left(\frac{L}{\varepsilon}+\frac{1}{\varepsilon^{2}}\right)$
REG [Mishchenko et al., 2020b] Independent $\mathcal{\widetilde{O}}\left(\frac{L}{\mu}+\frac{1}{\mu^{2}\varepsilon}\right)$ $\mathcal{O}\left(\frac{L}{\varepsilon}+\frac{1}{\varepsilon^{2}}\right)$
Extragradient [Carmon et al., 2019] Independent $\backslash$ $\mathcal{\widetilde{O}}\left(n+\frac{\sqrt{n}\overline{L}}{\varepsilon}\right)$
Mirror-prox [Carmon et al., 2019] Independent $\backslash$ $\mathcal{\widetilde{O}}\left(n+\frac{\sqrt{n}\overline{L}}{\varepsilon}\right)$
[…]
FoRB [Alacaoglu et al., 2021] Independent $\backslash$ $\mathcal{O}\left(n+\frac{n\overline{L}}{\varepsilon}\right)$
Extragradient [Alacaoglu and Malitsky, 2022] Independent $\mathcal{\widetilde{O}}\left(n+\frac{\sqrt{n}\overline{L}}{\mu}\right)$ $\mathcal{O}\left(n+\frac{\sqrt{n}\overline{L}}{\varepsilon}\right)$
Mirror-prox [Alacaoglu and Malitsky, 2022] Independent $\backslash$ $\mathcal{O}\left(n+\frac{\sqrt{n}\overline{L}}{\varepsilon}\right)$
[…]
  • Columns: Sampling = Deterministic, if considered non-stochastic method, Independent, if method uses independent choice of operator’s indices, RR / SO if method uses shuffling heuristic, Assumption = assumption on operator $F$, VR? = whether the method uses variance reduction technique.
  • Notation: $\mu$ = constant of strong monotonicity, $L$ = Lipschitz constant of $F$, $\overline{L}$ = Lipschitz in mean constant, i.e. $\nicefrac{{1}}{{n}}\sum\nolimits_{i=1}^{n}\|F_{i}(z_{1})-F_{i}(z_{2})\|\leqslant L\|z_{1}-z_{2}\|\penalty 10000\ \forall z_{1},z_{2}\in Z$, $n$ = size of the dataset, $\varepsilon$ = accuracy of the solution.
  • (1): This result is obtained with regularization trick: $\mu\sim\nicefrac{{\varepsilon}}{{D^{2}}}$.
Table 1: Comparison of the convergence results for the methods for solving VI.
In all these papers, the estimates were obtained through the formulation with an independent choice of the indices of the operator at each step of the algorithm. Regarding the shuffling heuristic, numerous studies explore methods suitable for addressing classic finite-sum minimization problems. In the work [Mishchenko et al., 2020a], the authors examined a classic SGD algorithm, and by introducing a new concept of variance specific to RR/SO, they matched the lower bounds in such scenarios. In the work [Malinovsky et al., 2023], the SVRG method with RR was considered. The authors actively utilized results from the work [Mishchenko et al., 2020a] and obtained improved rates. Additionally, there is a set of studies that considered methods incorporating the VR technique in the shuffling setup [Huang et al., 2021; Mokhtari et al., 2018; Ying et al., 2020]. However, there are currently no papers that employ the shuffling setting to solve variational inequalities. We aim to fill this gap.
Contributions. The main results can be summarized as follows.
$\bullet$ Considered algorithms. We consider two algorithms: Extragradient [Juditsky et al., 2011] and Extragradient that incorporate variance reduction [Alacaoglu and Malitsky, 2022], utilizing shuffling heuristics instead of independent index selection.
$\bullet$ Novel approach to proof. Since shuffling methods lack the property of unbiasedness of stochastic operators, it is essential to propose new approaches to demonstrate convergence. In this paper, we present a technique that enables us to "return" to the starting point of an epoch in which the property of unbiasedness is maintained.
$\bullet$ Convergence estimates. We provide the first theoretical convergence rates for shuffling methods applied to the finite-sum variational inequality problem. Our comprehensive analysis establishes upper bounds on convergence rates, extending the theoretical framework to encompass this important class of algorithms. In the case of Extragradient, our estimate on the linear term coincides with that for the method without shuffling. In the case of Extragradient with VR, we are the first to obtain a linear convergence estimate for methods with shuffling in the VI problem.
$\bullet$ Experiments. We conduct comprehensive experiments that emphasize the superiority of shuffling over the random index selection heuristic. We consider two classic practical applications: image denoising and adversarial training.
2  Setup
Assumptions. We present a list of assumptions under which we derive the main statements.
Assumption 1.
Each operator $F_{i}$ is $L$-Lipschitz, i.e., it satisfies $\|F_{i}(z_{1})-F_{i}(z_{2})\|\leq L\|z_{1}-z_{2}\|$ for any $z_{1},z_{2}\in Z$.
Assumption 2.
Each operator $F_{i}$ is $\mu$-strongly monotone, i.e., it satisfies $\langle F_{i}(z_{1})-F_{i}(z_{2}),z_{1}-z_{2}\rangle\geqslant\mu\|z_{1}-z_{2}\|^{2}$ for any $z_{1},z_{2}\in Z$.
Assumption 3.
Each stochastic operator $F_{i}$ and full operator $F$ is bounded at the point of the solution $z^{*}\in\text{dom}\penalty 10000\ g$, i.e. $\mathbb{E}\|F_{i}(z^{*})\|^{2}\leqslant\sigma_{*}^{2},\|F(z^{*})\|^{2}\leqslant\sigma_{*}^{2}$.
Proximal Algorithm. Earlier, we provided examples of the application of variational inequalities (Examples 1, 2). In many optimization problems, particularly in machine learning and signal processing, we often encounter the need to minimize a function of the same form, i.e., decompose it into two parts: a smooth differentiable function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and a possibly non-smooth function $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$. To solve this problem, we utilize the proximal gradient method. The core idea is to iteratively update the solution by combining Gradient descent on the smooth part $f$ and the proximal operator for the possibly non-smooth part $g$. We also assume that $g$ is proximal friendly, meaning the solution of the minimization problem on $g$ is achieved at minimal cost. The proximal operator of the function $g$ at a point $x$ is defined as:
$\text{prox}_{g}(z)=\arg\min_{y\in\mathbb{R}^{n}}\left\{g(y)+\frac{1}{2}\|y-z\|^{2}\right\},$
where $\|\cdot\|$ denotes the Euclidean norm. Using the proximal operator, the update step for solving the optimization problem can be expressed as
$z^{t+1}=\text{prox}_{\alpha_{t}g}\left(z^{t}-\alpha_{t}\nabla f(z^{t})\right).$
For us, the proximal operator plays a role, since (1) also uses a regularizer.
3  Algorithms and Convergence Analysis
3.1 Extragradient
The setting of shuffling lies in the fact that we do not choose the stochastic operator independently at each step of the method. Instead, we permute the sequence of indices and, at each iteration of the algorithm, we select the operator according to the new sequence. In this work, we focus on the Random Reshuffling and Shuffle Once techniques and provide appropriate Extragradient methods (Algorithms 3.1, 3.1).
RR Extragradient
1:Input: Starting point $z^{0}_{0}\in\mathbb{R}^{d}$
2:Parameter: Stepsize $\gamma$
3:for $s=0,1,2,\ldots,S-1$ do
4:  Generate a permutation $\pi_{0},\pi_{1},\ldots,\pi_{n-1}$ of sequence $\{1,2,\ldots,n\}$
5:  for $t=0,1,2,\ldots,n-1$ do
6:   $z^{t+\frac{1}{2}}_{s}=\text{prox}_{\gamma g}\left(z^{t}_{s}-\gamma F_{\pi_{s}^{t}}(z^{t}_{s})\right)$
7:   $z^{t+1}_{s}=\text{prox}_{\gamma g}\left(z^{t}_{s}-\gamma F_{\pi_{s}^{t}}(z^{t+\frac{1}{2}}_{s})\right)$
8:  end for
9:  $z_{s}^{n}=z_{s+1}^{0}$
10:end for
11:Output: $z^{n}_{S}$
SO Extragradient
1:Input: Starting point $z^{0}_{0}\in\mathbb{R}^{d}$
2:Parameter: Stepsize $\gamma$
3:Generate a permutation $\pi_{0},\pi_{1},\ldots,\pi_{n-1}$ of sequence $\{1,2,\ldots,n\}$
4:for $s=0,1,2,\ldots,S-1$ do
5:  for $t=0,1,2,\ldots,n-1$ do
6:   $z^{t+\frac{1}{2}}_{s}=\text{prox}_{\gamma g}\left(z^{t}_{s}-\gamma F_{\pi_{s}^{t}}(z^{t}_{s})\right)$
7:   $z^{t+1}_{s}=\text{prox}_{\gamma g}\left(z^{t}_{s}-\gamma F_{\pi_{s}^{t}}(z^{t+\frac{1}{2}}_{s})\right)$
8:  end for
9:  $z_{s}^{n}=z_{s+1}^{0}$
10:end for
11:Output: $z^{n}_{S}$
The analysis of shuffling methods has specific details. The key difference between shuffling and independent choice is that shuffling methods lack one essential feature: the unbiasedness of stochastic operators:
$\mathbb{E}_{\pi_{s}^{t}}\left[F_{\pi_{s}^{t}}(z_{s}^{t})\right]\neq\frac{1}{n}\sum\nolimits_{i=1}^{n}F_{\pi_{s}^{i}}(z_{s}^{t})=F(z_{s}^{t}).$
This restriction leads us to a more complex analysis and the use of non-standard techniques to prove the convergence of the shuffling methods. Nevertheless, at two points – $z_{s}^{0}$ and $z^{*}$, this equality holds. Indeed, the point $z_{s}^{0}$ is the initial point of the epoch, where we choose one random index from $n$. Additionally, the point $z^{*}$ does not depend on $t$. Thus, we can "go back" to the beginning of the epoch and leverage the unbiased operators. This technique is interesting not only in relation to shuffling methods. For example, it is applicable to methods that utilize Markov chains to select indices, as there is no unbiased property except at the correlation point of the chain. This is the key point of our analysis, and now, having established this, we present the main result of this section.
Theorem 1.
Suppose Assumptions 1, 2, 3 hold. Then for Algorithms 3.1, 3.1 with $\gamma\leqslant\min\left\{\frac{1}{2\mu n},\frac{1}{6L}\right\}$ after $S$ epochs,
$\|z_{S}^{n}-z^{*}\|^{2}\leqslant(1-\frac{\gamma\mu}{2})^{Sn}\|z^{0}_{0}-z^{*}\|^{2}+\frac{256\gamma n^{2}\sigma^{2}_{*}}{\mu}.$
Corollary 1.
Suppose Assumptions 1, 2, 3 hold. Then Algorithms 3.1, 3.1 with $\gamma\leqslant\min\left\{\frac{1}{2\mu n},\frac{1}{6L},\frac{2\log\left(\max\left\{2,\frac{\mu^{2}\|z_{0}^{0}-z^{*}\|^{2}T}{512n^{2}\sigma^{2}_{*}}\right\}\right)}{\mu T}\right\}$, to reach $\varepsilon$-accuracy, where $\varepsilon\sim\|z_{S}^{n}-z^{*}\|^{2}$, needs
$\mathcal{\widetilde{O}}\left(\left(n+\frac{L}{\mu}\right)\log\left(\frac{1}{\varepsilon}\right)+\frac{n^{2}\sigma^{2}_{*}}{\mu^{2}\varepsilon}\right)\penalty 10000\ \penalty 10000\ \text{iterations and oracle calls.}$
Remark 1.
We can transform the obtained estimation for the case of monotone stochastic operators. To achieve this, we utilize a regularization trick with $\mu\sim\frac{\varepsilon}{D}$. In particular, solving the problem with the operator $\hat{F}(z)=F(z)+\mu(z-z_{0}^{0})$ at accuracy $\frac{\varepsilon}{2}$ allows us to solve the problem (1) at accuracy $\varepsilon$, resulting in $\mathcal{\widetilde{O}}\left(n+\frac{L}{\varepsilon}+\frac{n^{2}}{\varepsilon^{3}}\right)$ for iteration and oracle complexity. This represents convergence in argument, which differs from the classic form.
Let us explain the result of the theorem. The form of the estimate is classic and appears in all stochastic methods for strongly convex minimization [Moulines and Bach, 2011; Stich, 2019] and strongly monotone VIs [Beznosikov et al., 2020; Mishchenko et al., 2020b]. We compare it with the results of related works. Our method is based on REG [Mishchenko et al., 2020b]. In this work, authors obtain $\mathcal{\widetilde{O}}\left(\frac{L}{\mu}+\frac{1}{\mu^{2}\varepsilon}\right)$ oracle complexity. Thus, our result represents a significant advancement in shuffling theory. Although there is a degradation in $n$ in the sublinear term, the estimation on the linear term coincides with that in the classic setting using an independent choice of stochastic operators. We also compare the result with the work [Juditsky et al., 2011]. The authors obtain $\mathcal{O}\left(\frac{L}{\sqrt{\varepsilon}}+\frac{1}{\varepsilon}\right)$. However, uniform bounds on the variance were required in this work, while we bound the variance only at the optimum. Note that, according to current theory, shuffling methods are no more effective than methods with independent sampling for classic minimization problems. Indeed, classic SGD rate in the non-convex case is $\mathcal{O}\left(\frac{L}{\sqrt{\varepsilon}}+\frac{L}{\varepsilon}\right)$. Therefore, prior works on shuffling in the minimization setup deliver $\mathcal{O}\left(\frac{nL}{\sqrt{\varepsilon}}+\frac{nL}{\varepsilon^{\nicefrac{{3}}{{4}}}}\right)$ [Mishchenko et al., 2020a; Mohtashami et al., 2022; Lu et al., 2022] and $\mathcal{O}\left(\frac{L}{\sqrt{\varepsilon}}+\frac{nL}{\varepsilon^{\nicefrac{{3}}{{4}}}}\right)$ [Koloskova et al., 2023] rates. Thus, despite the improvement in the asymptotics of the variance term, it deteriorates the theoretical estimate by a factor of $n$.
Let us focus on the second term in the estimation. In general, the sublinear term with $\sigma_{*}^{2}$ is not improved. However, for the finite-sum problem, this term can be eliminated by employing additional techniques, such as variance reduction.
3.2 Extragradient with Variance Reduction
RR/SO Extragradient with variance reduction
1:Input: Parameters: $z_{0}^{0},\omega_{0}^{0}$
2:Parameter: Stepsize $\gamma$, $\alpha\in(0,1)$
3:Generate a permutation $\pi_{0},\pi_{1},\ldots,\pi_{n-1}$ of sequence $\{1,2,\ldots,n\}$   // SO heuristic
4:for $s=0,1,\ldots$ do
5:  Generate a permutation $\pi_{0},\pi_{1},\ldots,\pi_{n-1}$ of sequence $\{1,2,\ldots,n\}$ // $R$ heuristic
6:  for $t=0,1,\ldots,n-1$ do
7:   $\overline{z}_{s}^{t}=\alpha z_{s}^{t}+(1-\alpha)\omega_{s}^{t}$
8:   $z_{s}^{t+\nicefrac{{1}}{{2}}}=\text{prox}_{\gamma g}\left(\overline{z}_{s}^{t}-\gamma F\left(\omega_{s}^{t}\right)\right)$
9:   $\hat{F}(z_{s}^{t+\nicefrac{{1}}{{2}}})=F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}})-F_{\pi_{s}^{t}}(\omega_{s}^{t})+F(\omega_{s}^{t})$
10:   $z_{s}^{t+1}=\text{prox}_{\gamma g}\left(\overline{z}_{s}^{t}-\gamma\hat{F}\left(z_{s}^{t+\nicefrac{{1}}{{2}}}\right)\right)$
11:    $\omega_{s}^{t+1}=\begin{cases}z_{s}^{t},\quad&\text{with probability}\quad p\\ \omega_{s}^{t}\quad&\text{with probability}\quad 1-p\end{cases}$
12:  end for
13:  $z_{s+1}^{0}=z_{s}^{n}$
14:  $\omega_{s+1}^{0}=\omega_{s}^{n}$
15:end for
16:Output: $z^{n}_{S}$
Now we utilize the variance reduction technique, which enhances the convergence of the algorithms by diminishing the impact of random fluctuations. This approach was not employed in the previously presented algorithms. We introduce a variant of the RR/SO Extragradient with a variance reduction algorithm (Algorithm 3.2) and provide the convergence results for this method. In the work [Malinovsky et al., 2023], where shuffling is investigated in variance reduction methods, the authors utilize a more classic version and compute $F(\omega_{s}^{t})$ at the beginning of each epoch. We consider an alternative and compute this full operator randomly with probability $p$. We set $p=\frac{1}{n}$, which implies that, on average, the full operator is updated once per epoch. Note that this choice does not increase the oracle complexity.
Theorem 2.
Suppose that Assumptions 1, 2 hold. Then for Algorithm 3.2 with $\gamma\leqslant\frac{(1-\alpha)\mu}{6L^{2}}$ and $p=\frac{1}{n}$ after $T$ iterations,
$V_{S}^{n}\leqslant\left(1-\frac{\gamma\mu}{4}\right)^{T}V_{0}^{0},$
where $V_{s}^{t}=\mathbb{E}\|z_{s}^{t}-z^{*}\|^{2}+\mathbb{E}\|\omega_{s}^{t}-z^{*}\|^{2}$.
Corollary 2.
Suppose that Assumptions 1, 2 hold. Then Algorithm 3.2 with $\gamma\leqslant\frac{(1-\alpha)\mu}{6L^{2}}$ and $p=\frac{1}{n}$, to reach $\varepsilon$-accuracy, where $\varepsilon\sim V_{S}^{n}$, needs
$\mathcal{O}\left(n\frac{L^{2}}{\mu^{2}}\log\left(\frac{1}{\varepsilon}\right)\right)\penalty 10000\ \penalty 10000\ \text{iterations and oracle calls,}$
where $V_{s}^{t}=\mathbb{E}\|z_{s}^{t}-z^{*}\|^{2}+\mathbb{E}\|\omega_{s}^{t}-z^{*}\|^{2}$.
Remark 2.
Similarly to Remark 1, we can apply our result in the monotone case through the regularization trick and obtain $\mathcal{\widetilde{O}}\bigl(n\frac{L^{2}}{\varepsilon^{2}}\bigr)$.
We remove the variance that arose in Theorem 1 and obtain linear convergence. Even though we obtain worse estimates than those in works that also use the variance reduction technique, such as [Alacaoglu and Malitsky, 2022; Alacaoglu et al., 2021; Chavdarova et al., 2019; Palaniappan and Bach, 2016] (see Table 1), there is a distinct explanation for this. According to current theory, methods with the shuffling heuristic are inferior to methods with independent sampling for the variance reduction methods [Malinovsky et al., 2023]. In the prior works, where the minimization of the strongly-convex objective with shuffling was considered, the following rates are obtained: $\mathcal{\widetilde{O}}\left(n^{2}\frac{L^{2}}{\mu^{2}}\right)$ [Gurbuzbalaban et al., 2017], $\mathcal{\widetilde{O}}\left(n\frac{L^{2}}{\mu^{2}}\right)$ [Ying et al., 2020], $\mathcal{\widetilde{O}}\left(n\frac{L^{\nicefrac{{3}}{{2}}}}{\mu^{\nicefrac{{3}}{{2}}}}\right)$ [Malinovsky et al., 2023]. At the same time, in classic minimization, variance reduction methods provide $\mathcal{\widetilde{O}}\left(n\frac{L}{\mu}\right)$ rates [Gorbunov et al., 2020]. Thus, it remains an open question whether it is possible to obtain theoretical convergence estimates for methods using shuffling heuristics that are equivalent to those for methods with independent index selection for the VI problem. Additionally, in the course of the work, no theoretical differences are revealed in the SO and RR techniques concerning the problem (1).
4  Experiments
In this section, we evaluate the proposed algorithms to demonstrate their practical applications by conducting experiments in two cases: image denoising and adversarial training.
4.1 Image Denoising
To formulate the image denoising problem [Chambolle and Pock, 2011], we consider the classic saddle point problem as presented in Example 2:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left[\langle Kx,y\rangle+G_{1}(x)-G_{2}(y)\right],$
where regularizers $G_{1}$ and $G_{2}$ are proper convex lower semicontinuous functions, and $K$ is a continuous linear operator. To proceed with image denoising, we consider $g$ as a given noisy image and $u$ as the solution we seek. We use the Cartesian grid with the step $h:\{(i\cdot h,j\cdot h)\}$. Thus, specifically for image denoising, we consider:
$\min_{u\in\mathcal{X}}\max_{p\in\mathcal{Y}}\left[\langle\nabla u,p\rangle_{\mathcal{Y}}+\nicefrac{{\lambda}}{{2}}\|u-g\|^{2}_{2}-\delta_{P}(p)\right],\vskip-5.69054pt$
where $p$ is a dual variable, $\delta_{P}(p)$ is the indicator function of the set $P$ defined as: $P=\{p\in\mathcal{Y}:\|p(x)\|\leq 1\}.$ The indicator function $\delta_{P}(p)$ is defined as zero if $p$ belongs to the set $P$, and infinity otherwise. We define operator $\nabla u$ as the difference between neighboring pixels in the grid, both horizontally and vertically, normalized by the step of the grid $h$. This formulation represents a saddle point problem, where we seek to minimize the first term with respect to $u$ while simultaneously maximizing the second term with respect to $p$. Using duality, we can write the final formulation of the problem as
$\min_{u\in\mathcal{X}}\max_{p\in\mathcal{Y}}\left[-\langle u,\text{div}\penalty 10000\ p\rangle_{\mathcal{X}}+\nicefrac{{\lambda}}{{2}}\|u-g\|^{2}_{2}-\delta_{P}(p)\right].$ (5)
To bring the problem to the form of a finite sum (4), we divide images into batches of equal squares. We consider two options: batches of size 4 and 8, according to the grid. Since the images are black and white, they are single-channel, which means that each batch is a square matrix with non-negative integers. It is also important to note that when calculating the gradient, the edges of the batch are processed according to the rule of adding a number equal to the nearest neighbor.
We compare RR/SO Extragradient with variance reduction (Algorithm 3.2) and Extragradient with variance reduction [Alacaoglu and Malitsky, 2022]. Analogously, we compare RR/SO Extragradient (Algorithms 3.1, 3.1) and Extragradient [Juditsky et al., 2011]. We select two images with different levels of additive zero-mean Gaussian noise: $\sigma=0.05$ and $\sigma=0.1$. Figures 1 and 2 provide a comparison of the proposed methods. Additional results for all considered methods on another image are presented in Figures 4, 5 in Appendix A.
Comparing the images, it is evident that algorithms incorporating shuffling perform better than those that do not, although the difference in the line graphs is subtle. While the results from the Independent Choice strategy appear sharper, they are also noisier compared to those from algorithms employing shuffling. Thus, utilizing shuffling techniques reduces noise more effectively.
Figure 1: Extragradient convergence on image with $\sigma=0.05$ on the problem (5).
Figure 2: Extragradient with VR convergence on image with $\sigma=0.05$ on the problem (5).
Hyperparameter details.
We configure the training with a fixed learning rate $\gamma=5\times 10^{-3}$ and a batch size of $8$. The probability parameter $p=\frac{1}{n}$ is determined based on the effective dataset size. However, for computational efficiency, we adapt it to the mini-batch context. Training is conducted with random state $50$.
4.2 Adversarial Training
Next, we address an adversarial training problem. We formulate it in the following manner:
$\min_{w\in\mathbb{R}^{d}}\max_{\|r_{i}\|\leqslant D}\left[\frac{1}{2N}\sum\limits_{i=1}^{N}\left(w^{T}\left(x_{i}+r_{i}\right)-y_{i}\right)^{2}+\frac{\lambda}{2}\|w\|^{2}-\frac{\beta}{2}\|r\|^{2}\right],$ (6)
where the samples correspond to features $x_{i}$ and targets $y_{i}$. We evaluate this issue across several datasets: mushrooms, a9a, and w8a, sourced from the LIBSVM library [Chang and Lin, 2011]. A brief description of these datasets is provided in Table 2, Appendix A. The results are presented in Figure 3.
(a) mushrooms
(b) a9a
(c) w8a
Figure 3: Extragradient with and without VR compared using various shuffling heuristics on the datasets shown above for the problem (6).
As shown in the plots, Extragradient and Extragradient with the VR algorithms using the independent choice of indices demonstrate worse performance compared to those using shuffling methods. In this series of experiments, RR exhibits better performance than other shuffling methods and significantly outperforms the non-shuffled versions.
Hyperparameter Details. We configure the training with a fixed learning rate $\gamma=0.01$ and a batch size of $4,4,16$ on mushrooms, a9a and w8a datasets, respectively. The probability parameter $p=\frac{1}{n}$ is determined based on the effective dataset size; however, for computational efficiency, we adapt it to the mini-batch context. Training is conducted with random state $50$.
5  Conclusion
In this paper, we consider stochastic algorithms for solving the problem of variational inequalities. Specifically, we explore the influence of shuffling heuristics on stochastic methods. For the first time, a theoretical analysis of two versions of the Extragradient algorithm, both with and without variance reduction involving shuffling, is presented. Empirical results on image denoising and adversarial training tasks confirm the applicability of the methods in practice. Nevertheless, the convergence rate of the Extragradient algorithm with variance reduction is not optimal (see Corollary 2). Therefore, its improvement can be considered as a future work.
Appendix Supplementary Materials for Shuffling Heuristic in Variational Inequalities: Establishing New Convergence Guarantees
Appendix A Additional Experiments
In this section, we present additional experiments that have been performed (Figures 4, 5). Similar to the previous experiments, we observe a consistent pattern: methods incorporating shuffling techniques outperform those without shuffling. These results further confirm the effectiveness of shuffling techniques in addressing the denoising problem on another image with higher $\sigma$.
Figure 4: Extragradient convergence on image with $\sigma=0.1$ on the problem (5).
Name
Number of
Instances
Number of
Features
Number of
Classes
mushrooms 8,124 112 2
a9a 32,561 123 2
w8a 49,749 300 2
Table 2: Summary of Datasets
Figure 5: Extragradient with VR convergence on image with $\sigma=0.1$ on the problem (5).
The datasets used for the experiments on the adversarial training include mushrooms, a9a, and w8a. These datasets vary in size and complexity (see Table 2 for details), providing a comprehensive evaluation of our proposed algorithms in the context of adversarial training.
Appendix B Basic Inequalities
For all vectors $x,y,\{x_{i}\}_{i=1}^{n}$ in $\mathbb{R}^{d}$ with a positive scalar $\alpha$, the following holds:
$\langle x,y\rangle\penalty 10000\$$\leqslant\penalty 10000\ \frac{\|x\|^{2}}{2\alpha}+\frac{\alpha\|y\|^{2}}{2},$ (Scalar)
$2\langle x,y\rangle\penalty 10000\$$=\penalty 10000\ \|x+y\|^{2}-\|x\|^{2}-\|y\|^{2},$ (Norm)
$-2\|x\|^{2}\penalty 10000\$$\leqslant\penalty 10000\ -\|x+y\|^{2}+2\|y\|^{2},$ (CS)
$\left\|\sum_{i=1}^{n}x_{i}\right\|^{2}$ $\leqslant\penalty 10000\ n\sum_{i=1}^{n}\|x_{i}\|^{2},$ (Sum)
$\left\|\text{prox}_{\gamma g}(x)-\text{prox}_{\gamma g}(y)\right\|^{2}$ $\leqslant\left\|x-y\right\|^{2}.$ (Prox)
Appendix C Extragradient
Theorem 1. Suppose Assumptions 1, 2 hold. Then for Algorithms 3.1, 3.1 with $\gamma\leqslant\min\left\{\frac{1}{2\mu n},\frac{1}{6L}\right\}$ after $S$ epochs,
$\|z_{S}^{n}-z^{*}\|^{2}\leqslant(1-\frac{\gamma\mu}{2})^{Sn}\|z^{0}_{0}-z^{*}\|^{2}+\frac{256\gamma n^{2}\sigma^{2}_{*}}{\mu}.$
Proof.
We begin with the standard prox-inequality:
$\hat{z}=\text{prox}_{g}(z)\Longleftrightarrow\langle\hat{z}-z,u-\hat{z}\rangle\geqslant g(\hat{z})-g(u),\quad\forall u\in\mathcal{Z}.$ (7)
Substituting both steps of Algorithm 3.1 (Algorithm 3.1) into (7), we derive:
$\langle z_{s}^{t+1}-z_{s}^{t}+\gamma F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z^{*}-z_{s}^{t+1}\rangle$ $\geqslant\gamma(g(z_{s}^{t+1})-g(z^{*})),$
$\langle z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}+\gamma F_{\pi_{s}^{t}}(z_{s}^{t}),z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$ $\geqslant\gamma(g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z_{s}^{t+1})).$
Summing the inequalities, we obtain:
$\gamma(g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z^{*}))$ $\leqslant\langle z_{s}^{t+1}-z_{s}^{t},z^{*}-z_{s}^{t+1}\rangle+\langle z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t},z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$
$\quad+\gamma\langle F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z^{*}-z_{s}^{t+1}\rangle+\gamma\langle F_{\pi_{s}^{t}}(z_{s}^{t}),z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle.$
Now, we add and subtract $z_{s}^{t+\nicefrac{{1}}{{2}}}$ to the right part of the third scalar product. Thus, rearranging terms, we arrive at
$\gamma(g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z^{*}))$ $\leqslant\langle z_{s}^{t+1}-z_{s}^{t},z^{*}-z_{s}^{t+1}\rangle+\langle z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t},z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$
$\quad+\gamma\langle F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z^{*}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$
$\quad+\gamma\langle F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}})-F_{\pi_{s}^{t}}(z_{s}^{t}),z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t+1}\rangle.$ (8)
We want to rewrite the first two scalar products. We use (Norm). Thus, we arrive at
$2\langle z_{s}^{t+1}-z_{s}^{t},z^{*}-z_{s}^{t+1}\rangle$ $=\|z_{s}^{t}-z^{*}\|^{2}-\|z_{s}^{t+1}-z_{s}^{t}\|^{2}-\|z^{*}-z_{s}^{t+1}\|^{2},$
$2\langle z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t},z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$ $=\|z_{s}^{t+1}-z_{s}^{t}\|^{2}-\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}\|^{2}-\|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2}.$
Substituting this into (8), we obtain:
$\|z_{s}^{t+1}-z^{*}\|^{2}$ $\leqslant\|z_{s}^{t}-z^{*}\|^{2}-2\gamma\left(\langle F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\rangle+g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z^{*})\right)$
$\quad+2\gamma\langle F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}})-F_{\pi_{s}^{t}}(z_{s}^{t}),z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t+1}\rangle$
$\quad-\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}\|^{2}-\|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2}.$
Now, applying (Scalar) and Assumption 1 to the second scalar product, we obtain:
$\|z_{s}^{t+1}-z^{*}\|^{2}$ $\penalty 10000\ \penalty 10000\ \leqslant\penalty 10000\ \|z_{s}^{t}-z^{*}\|^{2}-2\gamma\bigl(\langle F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\rangle$
$\penalty 10000\ \penalty 10000\ \quad+\penalty 10000\ g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z^{*})\bigr)+\gamma^{2}\|F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}})-F_{\pi_{s}^{t}}(z_{s}^{t})\|^{2}$
$\penalty 10000\ \penalty 10000\ \quad+\penalty 10000\ \|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2}-\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}\|^{2}-\|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2}$
[…]
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -2\gamma\left(\underbrace{\langle F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\rangle+g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z^{*})}_{T_{1}}\right).$ (9)
To estimate the $T_{1}$ term, we compute the expectation:
$\mathbb{E}T_{1}$ $\penalty 10000\ \penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathbb{E}\langle F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\rangle+g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z^{*})$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathbb{E}\langle F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}})-F_{\pi_{s}^{t}}(z^{*}),z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\rangle$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \quad\quad+\penalty 10000\ \mathbb{E}\langle F_{\pi_{s}^{t}}(z^{*}),z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\rangle+g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z^{*})$
[…]
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \quad\quad+\penalty 10000\ \mathbb{E}\langle F_{\pi_{s}^{t}}(z^{*}),z_{s}^{0}-z^{*}\rangle+g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z^{*}).$
Now we focus on the second scalar product. Utilizing the tower property and the unbiasedness of the stochastic operator at the points $z_{s}^{0}$ and $z^{*}$, we obtain $\mathbb{E}\left[\mathbb{E}_{t}\left[F_{\pi_{s}^{t}}(z^{*})|z_{s}^{0}-z^{*}\right]\right]=F(z^{*})$. Thus, we continue the estimation of $\mathbb{E}T_{1}$:
[…]
$\penalty 10000\ \penalty 10000\ \quad\penalty 10000\ \penalty 10000\ \penalty 10000\ +\langle F(z^{*}),z_{s}^{0}-z^{*}\rangle+g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z^{*})$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ =\quad\mu\mathbb{E}\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\|^{2}-\frac{\gamma}{2\beta}\mathbb{E}\|F_{\pi_{s}^{t}}(z^{*})\|^{2}-\frac{\beta}{2\gamma}\mathbb{E}\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{0}\|^{2}$
[…]
$\penalty 10000\ \penalty 10000\ \quad\penalty 10000\ \penalty 10000\ \penalty 10000\ -\frac{\gamma}{2\beta}\|F(z^{*})\|^{2}-\frac{\beta}{\gamma}\mathbb{E}\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{0}\|^{2}.$
Here we introduce $\beta>0$, which we will define later. Substituting this inequality into (9) yields:
$\mathbb{E}\|z_{s}^{t+1}-z^{*}\|^{2}$ $\penalty 10000\ \penalty 10000\ \quad\leqslant\quad\penalty 10000\ \mathbb{E}\|z_{s}^{t}-z^{*}\|^{2}-2\gamma\mu\mathbb{E}\|z_{s}^{t+\frac{1}{2}}-z^{*}\|^{2}$
$\quad\quad\quad\penalty 10000\ \penalty 10000\ +(\gamma^{2}L^{2}-1)\mathbb{E}\|z_{s}^{t+\frac{1}{2}}-z_{s}^{t}\|^{2}+\frac{\gamma^{2}}{\beta}\mathbb{E}\|F_{\pi_{s}^{t}}(z^{*})\|^{2}$
$\quad\quad\quad\penalty 10000\ \penalty 10000\ +\frac{\gamma^{2}}{\beta}\|F(z^{*})\|^{2}+2\beta\mathbb{E}\|z_{s}^{t+\frac{1}{2}}-z_{s}^{0}\|^{2}$
[…]
$\quad\quad\quad\penalty 10000\ \penalty 10000\ +(\gamma^{2}L^{2}+2\gamma\mu-1)\mathbb{E}\|z_{s}^{t+\frac{1}{2}}-z_{s}^{t}\|^{2}.$ (10)
Now, we evaluate the $\|z_{s}^{t+\frac{1}{2}}-z_{s}^{0}\|^{2}$ term:
$\|z_{s}^{t+\frac{1}{2}}-z_{s}^{0}\|^{2}$ $\penalty 10000\ \penalty 10000\ \leqslant\left(1+\frac{1}{a}\right)\|z_{s}^{t+\frac{1}{2}}-z_{s}^{t}\|^{2}+(1+a)\|z_{s}^{t}-z_{s}^{0}\|^{2}$
$\penalty 10000\ \penalty 10000\ \leqslant\left(1+\frac{1}{a}\right)\|z_{s}^{t+\frac{1}{2}}-z_{s}^{t}\|^{2}+(1+a)\left(1+\frac{1}{b}\right)\|z_{s}^{t}-z_{s}^{t-\frac{1}{2}}\|^{2}$
$\penalty 10000\ \penalty 10000\ \quad+(1+a)(1+b)\|z_{s}^{t-\frac{1}{2}}-z_{s}^{0}\|^{2}$
$\penalty 10000\ \penalty 10000\ =\left(1+\frac{1}{a}\right)\|z_{s}^{t+\frac{1}{2}}-z_{s}^{t}\|^{2}$
$\penalty 10000\ \penalty 10000\ \quad+(1+a)\left(1+\frac{1}{b}\right)\Bigl\|\text{prox}_{\gamma g}\left(z_{s}^{t-1}-\gamma F_{\pi_{s}^{t}}(z_{s}^{t-\frac{1}{2}})\right)$
$\penalty 10000\ \penalty 10000\ \quad-\text{prox}_{\gamma g}\left(z_{s}^{t-1}-\gamma F_{\pi_{s}^{t}}(z_{s}^{t-1})\right)\Bigr\|^{2}+(1+a)(1+b)\|z_{s}^{t-\frac{1}{2}}-z_{s}^{0}\|^{2}$
[…]
$\penalty 10000\ \penalty 10000\ \quad+(1+a)\left(1+\frac{1}{b}\right)\left\|\gamma F_{\pi_{s}^{t}}(z_{s}^{t-\frac{1}{2}})-\gamma F_{\pi_{s}^{t}}(z_{s}^{t-1})\right\|^{2}$
$\penalty 10000\ \penalty 10000\ \quad+(1+a)(1+b)\|z_{s}^{t-\frac{1}{2}}-z_{s}^{0}\|^{2}$
$\penalty 10000\ \penalty 10000\ \leqslant\left(1+\frac{1}{a}\right)\|z_{s}^{t+\frac{1}{2}}-z_{s}^{t}\|^{2}$
$\penalty 10000\ \penalty 10000\ \quad+(1+a)\left(1+\frac{1}{b}\right)\gamma^{2}L^{2}\|z_{s}^{t-\frac{1}{2}}-z_{s}^{t-1}\|^{2}$
$\penalty 10000\ \penalty 10000\ \quad+(1+a)(1+b)\|z_{s}^{t-\frac{1}{2}}-z_{s}^{0}\|^{2}$
$\penalty 10000\ \penalty 10000\ \leqslant\left(1+\frac{1}{a}\right)\|z_{s}^{t+\frac{1}{2}}-z_{s}^{t}\|^{2}+\sum\limits_{i=1}^{t-1}\|z_{s}^{i+\frac{1}{2}}-z_{s}^{i}\|^{2}$
$\penalty 10000\ \penalty 10000\ \quad\penalty 10000\ \cdot\left(\left(1+\frac{1}{a}\right)(1+a)(1+b)+(1+a)\left(1+\frac{1}{b}\right)\gamma^{2}L^{2}\right)$
$\penalty 10000\ \penalty 10000\ \quad\penalty 10000\ \cdot\left[(1+a)(1+b)\right]^{t-1-i}$
$\penalty 10000\ \penalty 10000\ \quad+\left((1+a)\left(1+\frac{1}{b}\right)+1\right)\left[(1+a)(1+b)\right]^{t}\|z_{s}^{\frac{1}{2}}-z_{s}^{0}\|^{2}.$ (11)
We choose $a=b=\frac{1}{n}$ and consider the coefficients for all three terms.
$1+\frac{1}{a}$ $=1+n,$
$\left(\left(1+\frac{1}{a}\right)+\frac{1}{b}\gamma^{2}L^{2}\right)\left[(1+a)(1+b)\right]^{t-i}$ $=(1+n+n\gamma^{2}L^{2})\left(1+\frac{1}{n}\right)^{2(t-i)},$
$\left((1+a)\left(1+\frac{1}{b}\right)+1\right)\left[(1+a)(1+b)\right]$ $^{t-i}\bigg|_{i=0}$
$=\left(3+\frac{1}{n}+n\right)\left(1+\frac{1}{n}\right)^{2(t-i)}\bigg|_{i=0}.$
We can evaluate the smaller terms from above by the largest one and consolidate them into a single sum. Thus, (11) transforms to
$\|z_{s}^{t+\frac{1}{2}}-z_{s}^{0}\|^{2}\leqslant\sum\limits_{i=0}^{t}\|z_{s}^{i+\frac{1}{2}}-z_{s}^{i}\|^{2}\left(1+n+n\gamma^{2}L^{2}\right)\left(1+\frac{1}{n}\right)^{2(t-i)}.$
Let us substitute the obtained inequality into (10):
$\mathbb{E}\|z_{s}^{t+1}-z^{*}\|^{2}$ $\leqslant(1-\gamma\mu)\mathbb{E}\|z_{s}^{t}-z^{*}\|^{2}+(\gamma^{2}L^{2}+2\gamma\mu-1)\mathbb{E}\|z_{s}^{t+\frac{1}{2}}-z_{s}^{t}\|^{2}\penalty 10000\$
$\quad+\frac{2\gamma^{2}}{\beta}\sigma_{*}^{2}+2\beta\sum\limits_{i=0}^{t}\mathbb{E}\|z_{s}^{i+\frac{1}{2}}-z_{s}^{i}\|^{2}\left(1+n+n\gamma^{2}L^{2}\right)\left(1+\frac{1}{n}\right)^{2(t-i)}.$ (12)
Now we define a new sequence that contains iteration points in all epochs:
$\widetilde{z}_{k}=z_{t+sn}.$
Thus, additionally considering $\left(1+\frac{1}{n}\right)^{2(t-i)}\leqslant\left(1+\frac{1}{n}\right)^{2n}\leqslant e^{2}\leqslant 8$, we can rewrite (12) in the following form:
$\mathbb{E}\|\widetilde{z}_{k+1}-z^{*}\|^{2}$ $\leqslant(1-\gamma\mu)\mathbb{E}\|\widetilde{z}_{k}-z^{*}\|^{2}+(\gamma^{2}L^{2}+2\gamma\mu-1)\mathbb{E}\|\widetilde{z}_{k+\frac{1}{2}}-\widetilde{z}_{k}\|^{2}\penalty 10000\$
$\quad+\frac{2\gamma^{2}}{\beta}\sigma_{*}^{2}+16\beta\sum\limits_{i=0}^{n}\mathbb{E}\|\widetilde{z}_{k-i+\frac{1}{2}}-\widetilde{z}_{k-i}\|^{2}\left(1+n+n\gamma^{2}L^{2}\right).$
Let us pay attention to the $\sum\limits_{i=0}^{n}\mathbb{E}\|\widetilde{z}_{k-i+\frac{1}{2}}-\widetilde{z}_{k-i}\|^{2}$ term in the obtained inequality. For the original sequence, this term represented the sum of the norms from the beginning of the current epoch to the current iteration $t$ and could contain a maximum of $n$ terms. Thus, a new expression that includes the sum of $n$ norms up to the current iteration $k$ serves as an upper bound, confirming that our expression is correct. Now we define $p_{k}=p^{k}=\left(1-\frac{\gamma\mu}{2}\right)^{-k}$ and summarize both sides over all iterations with coefficients $p_{k}$:
$\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k+1}-z^{*}\|^{2}$ $\leqslant(1-\gamma\mu)\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k}-z^{*}\|^{2}$
$\quad+(\gamma^{2}L^{2}+2\gamma\mu-1)\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k+\frac{1}{2}}-\widetilde{z}_{k}\|^{2}$
$\quad+\sum\limits_{k=0}^{Sn-1}p_{k}\sum\limits_{i=0}^{n}\mathbb{E}\|\widetilde{z}_{k-i+\frac{1}{2}}-\widetilde{z}_{k-i}\|^{2}$
$\quad\cdot 16\beta\left(1+n+n\gamma^{2}L^{2}\right)+\frac{2\gamma^{2}\sigma_{*}^{2}}{\beta}\sum\limits_{k=0}^{Sn-1}p_{k}.$ (13)
Now we need to estimate the following term:
$\sum\limits_{k=0}^{Sn-1}p_{k}\sum\limits_{i=0}^{n}\mathbb{E}\|\widetilde{z}_{k-i+\frac{1}{2}}-\widetilde{z}_{k-i}\|^{2}$ $\leqslant p_{n}\sum\limits_{k=0}^{Sn-1}\sum\limits_{i=0}^{n}p_{k-i}\mathbb{E}\|\widetilde{z}_{k-i+\frac{1}{2}}-\widetilde{z}_{k-i}\|^{2}$
$\leqslant p_{n}n\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k+\frac{1}{2}}-\widetilde{z}_{k}\|^{2}.$
Note that we define points $\widetilde{z}_{-n},\widetilde{z}_{-n+\frac{1}{2}},\ldots,\widetilde{z}_{-\frac{1}{2}}$ by shifting the sequence $\{\widetilde{z}_{k}\}$ on $n$ points. Since $p_{n}=\left(1-\frac{\gamma\mu}{2}\right)^{-n}=\left(1-\frac{\gamma\mu n}{2n}\right)^{-n}\leqslant e^{\frac{\gamma\mu n}{2}}$, we choose $\gamma\leqslant\frac{1}{2\mu n}$ and obtain $p_{n}\leqslant e^{\frac{1}{4}}\leqslant 2$. Substituting this into (13), we obtain:
$\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k+1}-z^{*}\|^{2}$ $\leqslant(1-\gamma\mu)\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k}-z^{*}\|^{2}+\frac{2\gamma^{2}\sigma_{*}^{2}}{\beta}\sum\limits_{k=0}^{Sn-1}p_{k}$
$\quad+(\gamma^{2}L^{2}+2\gamma\mu-1)\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k+\frac{1}{2}}-\widetilde{z}_{k}\|^{2}$
$\quad+32\beta\left(1+n+n\gamma^{2}L^{2}\right)n\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k+\frac{1}{2}}-\widetilde{z}_{k}\|^{2}.$
We consider the coefficient before $\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k+\frac{1}{2}}-\widetilde{z}_{k}\|^{2}$ and we make it negative by selecting $\gamma$ and $\beta$.
$32\beta(1+n+n\gamma^{2}L^{2})n+\gamma^{2}L^{2}+2\gamma\mu-1\leqslant 0$
We need $\gamma\leqslant\frac{1}{6L},\beta=\frac{1}{64n^{2}}.$ Then to satisfy the previous estimate on gamma we finally put $\gamma\leqslant\min\left\{\frac{1}{2\mu n},\frac{1}{6L}\right\}$ and, assuming $n>3$, have
$\frac{1}{2n}+\frac{1}{2}+\frac{1}{72}+\frac{1}{36}+\frac{1}{3}-1\leqslant 0.$
In this way,
$\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k+1}-z^{*}\|^{2}\leqslant(1-\gamma\mu)\sum\limits_{k=0}^{Sn-1}p_{k}\mathbb{E}\|\widetilde{z}_{k}-z^{*}\|^{2}+\frac{2\gamma^{2}\sigma_{*}^{2}}{\beta}\sum\limits_{k=0}^{Sn-1}p_{k}.$
Thus, substituting definition of $p_{t}$, we obtain:
$\sum\limits_{k=0}^{Sn-1}\left(1-\frac{\gamma\mu}{2}\right)^{-k}\mathbb{E}\|\widetilde{z}_{k+1}-z^{*}\|^{2}$ $\leqslant\sum\limits_{k=0}^{Sn-1}\left(1-\frac{\gamma\mu}{2}\right)^{-k+1}\mathbb{E}\|\widetilde{z}_{k}-z^{*}\|^{2}$
$\quad+\frac{2\gamma^{2}\sigma_{*}^{2}}{\beta}\sum\limits_{k=0}^{Sn-1}\left(1-\frac{\gamma\mu}{2}\right)^{-k},$
$\left(1-\frac{\gamma\mu}{2}\right)^{-(Sn-1)}\mathbb{E}\|\widetilde{z}_{Sn}-z^{*}\|^{2}$ $\leqslant\left(1-\frac{\gamma\mu}{2}\right)\mathbb{E}\|\widetilde{z}_{0}-z^{*}\|^{2}$
$\quad+\frac{2\gamma^{2}\sigma_{*}^{2}}{\beta}\sum\limits_{k=0}^{Sn-1}\left(1-\frac{\gamma\mu}{2}\right)^{-k},$
$\mathbb{E}\|\widetilde{z}_{Sn}-z^{*}\|^{2}$ $\leqslant\left(1-\frac{\gamma\mu}{2}\right)^{Sn}\mathbb{E}\|\widetilde{z}_{0}-z^{*}\|^{2}$
$\quad+\frac{2\gamma^{2}\sigma_{*}^{2}}{\beta}\sum\limits_{k=0}^{Sn-1}\left(1-\frac{\gamma\mu}{2}\right)^{Sn-k-1}$
$=\left(1-\frac{\gamma\mu}{2}\right)^{Sn}\mathbb{E}\|\widetilde{z}_{0}-z^{*}\|^{2}$
$\quad+\frac{2\gamma^{2}\sigma_{*}^{2}}{\beta}\sum\limits_{k=0}^{Sn-1}\left(1-\frac{\gamma\mu}{2}\right)^{k}.$
Finally, estimating the geometric progression in the last term as $\sum\limits_{k=0}^{Sn-1}\left(1-\frac{\gamma\mu}{2}\right)^{k}\leqslant\frac{2}{\gamma\mu}$, we can write the final statement of the theorem:
$\mathbb{E}\|z_{S}^{n}-z^{*}\|^{2}$ $\leqslant\left(1-\frac{\gamma\mu}{2}\right)^{Sn}\mathbb{E}\|z_{0}^{0}-z^{*}\|^{2}+\frac{4\gamma\sigma_{*}^{2}}{\beta\mu}$
$=\left(1-\frac{\gamma\mu}{2}\right)^{Sn}\mathbb{E}\|z_{0}^{0}-z^{*}\|^{2}+\frac{256\gamma n^{2}\sigma_{*}^{2}}{\mu}.$
Corollary 1. Suppose Assumptions 1, 2 hold. Then Algorithms 3.1, 3.1 with
$\gamma\leqslant\min\left\{\frac{1}{2\mu n},\frac{1}{6L},\frac{2\log\left(\max\left\{2,\frac{\mu^{2}\|z_{0}^{0}-z^{*}\|^{2}T}{512n^{2}\sigma^{2}_{*}}\right\}\right)}{\mu T}\right\}$, to reach $\varepsilon$-accuracy, where $\varepsilon\sim\|z_{S}^{n}-z^{*}\|^{2}$, needs
$\mathcal{\widetilde{O}}\left(\left(n+\frac{L}{\mu}\right)\log\left(\frac{1}{\varepsilon}\right)+\frac{n^{2}\sigma^{2}_{*}}{\mu^{2}\varepsilon}\right)\penalty 10000\ \penalty 10000\ \textit{iterations and oracle calls.}$
Proof.
For the result obtained in Theorem 1, we utilize Lemma 2 from [Stich, 2019] and, using special tuning of $\gamma$, such as $\gamma\leqslant\min\left\{\frac{1}{2\mu n},\frac{1}{6L},\frac{2\log\left(\max\left\{2,\frac{\mu^{2}\|z_{0}^{0}-z^{*}\|^{2}T}{512n^{2}\sigma^{2}_{*}}\right\}\right)}{\mu T}\right\}$, we obtain that we need $\mathcal{\widetilde{O}}\left(\left(n+\frac{L}{\mu}\right)\log\left(\frac{1}{\varepsilon}\right)+\frac{n^{2}\sigma^{2}_{*}}{\mu^{2}\varepsilon}\right)$ iterations and oracle calls to reach $\varepsilon$-accuracy, where $\varepsilon\sim\|z_{S}^{n}-z^{*}\|^{2}$. ∎
Appendix D Extragradient with Variance Reduction
Theorem 2. Suppose that Assumptions 1, 2 hold. Then for Algorithm 3.2 with $\gamma\leqslant\frac{(1-\alpha)\mu}{6L^{2}},p=\frac{1}{n}$ and $V_{s}^{t}=\mathbb{E}\|z_{s}^{t}-z^{*}\|^{2}+\mathbb{E}\|\omega_{s}^{t}-z^{*}\|^{2}$, after $T$ iterations we have
$V_{S}^{n}\leqslant\left(1-\frac{\gamma\mu}{4}\right)^{T}V_{0}^{0}.$
Proof.
We start with substituting both steps of Algorithm 3.2 to (7):
$\langle z_{s}^{t+1}-\overline{z}_{s}^{t}+\gamma\hat{F}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z^{*}-z_{s}^{t+1}\rangle$ $\geqslant\gamma(g(z_{s}^{t+1})-g(z^{*})),$
$\langle z_{s}^{t+\nicefrac{{1}}{{2}}}-\overline{z}_{s}^{t}+\gamma F(\omega_{s}^{t}),z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$ $\geqslant\gamma(g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z_{s}^{t+1})).$
Let us summarize this two inequalities:
$\gamma(g(z_{s}^{t+\nicefrac{{1}}{{2}}})-g(z^{*}))$ $\leqslant\langle z_{s}^{t+1}-\overline{z}_{s}^{t},z^{*}-z_{s}^{t+1}\rangle+\langle z_{s}^{t+\nicefrac{{1}}{{2}}}-\overline{z}_{s}^{t},z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$
$\quad+\gamma\langle\hat{F}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z^{*}-z_{s}^{t+1}\rangle+\gamma\langle F(\omega_{s}^{t}),z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle.$
Now, we add and subtract $z_{s}^{t+\nicefrac{{1}}{{2}}}$ to the right part of the third scalar product. Thus, rearranging terms and utilizing the definition of $\hat{F}(z_{s}^{t+\nicefrac{{1}}{{2}}})$, we arrive at:
$\underbrace{\langle z_{s}^{t+1}-\overline{z}_{s}^{t},z^{*}-z_{s}^{t+1}\rangle}_{T_{1}}+\underbrace{\langle z_{s}^{t+\nicefrac{{1}}{{2}}}-\overline{z}_{s}^{t},z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle}_{T_{2}}$
$\quad+\underbrace{\gamma\langle F_{\pi_{s}^{t}}(\omega_{s}^{t})-F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle}_{T_{3}}$
$\quad+\underbrace{\gamma\langle\hat{F}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z^{*}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle+\gamma(g(z^{*})-g(z_{s}^{t+\nicefrac{{1}}{{2}}}))}_{T_{4}}\geqslant 0.$ (14)
We defined terms as $T_{1},T_{2},T_{3},T_{4}$, respectively. Let us estimate them separately. We start with $T_{1}$ and $T_{2}$. To estimate them, firstly, we use the definition of $\overline{z}_{s}^{t}$ and, secondly, use (Norm). Thus, we obtain:
$2T_{1}$ $=2\langle z_{s}^{t+1}-\overline{z}_{s}^{t},z^{*}-z_{s}^{t+1}\rangle$
$=2\alpha\langle z_{s}^{t+1}-z_{s}^{t},z^{*}-z_{s}^{t+1}\rangle+2(1-\alpha)\langle z_{s}^{t+1}-\omega_{s}^{t},z^{*}-z_{s}^{t+1}\rangle$
$=\alpha(\|z^{*}-z_{s}^{t}\|^{2}-\|z_{s}^{t+1}-z_{s}^{t}\|^{2}-\|z^{*}-z_{s}^{t+1}\|^{2})$
$\quad+(1-\alpha)(\|z^{*}-\omega_{s}^{t}\|^{2}-\|z_{s}^{t+1}-\omega_{s}^{t}\|^{2}-\|z^{*}-z_{s}^{t+1}\|^{2})$
$=\alpha\|z_{s}^{t}-z^{*}\|^{2}-\|z_{s}^{t+1}-z^{*}\|^{2}+(1-\alpha)\|\omega_{s}^{t}-z^{*}\|^{2}$
$\quad-\alpha\|z_{s}^{t+1}-z_{s}^{t}\|^{2}-(1-\alpha)\|z_{s}^{t+1}-\omega_{s}^{t}\|^{2}.$
The same holds for $T_{2}$:
$2T_{2}$ $=2\langle z_{s}^{t+\nicefrac{{1}}{{2}}}-\overline{z}_{s}^{t},z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$
$=2\alpha\langle z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t},z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle+2(1-\alpha)\langle z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t},z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$
$=\alpha(\|z_{s}^{t+1}-z_{s}^{t}\|^{2}-\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}\|^{2}-\|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2})$
$\quad+(1-\alpha)(\|z_{s}^{t+1}-\omega_{s}^{t}\|^{2}-\|z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t}\|^{2}-\|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2})$
$=\alpha\|z_{s}^{t+1}-z_{s}^{t}\|^{2}-\|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2}+(1-\alpha)\|z_{s}^{t+1}-\omega_{s}^{t}\|^{2}$
$\quad-\alpha\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}\|^{2}-(1-\alpha)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t}\|^{2}.$
Now, we moving to the estimate of $T_{3}$:
$2T_{3}$ = $\penalty 10000\ \penalty 10000\ \penalty 10000\ =\penalty 10000\ \penalty 10000\ \penalty 10000\ 2\gamma\langle F_{\pi_{s}^{t}}(\omega_{s}^{t})-F_{\pi_{s}^{t}}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$
[…]
Here we introduced $\tau>0$, which we will define later. Last, we do the same for $T_{4}$:
$2T_{4}$ $\penalty 10000\ \quad\quad=\quad\quad 2\gamma\langle\hat{F}(z_{s}^{t+\nicefrac{{1}}{{2}}}),z^{*}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle+2\gamma(g(z^{*})-g(z_{s}^{t+\nicefrac{{1}}{{2}}}))$
$\penalty 10000\ \quad\quad=\quad\quad 2\gamma\langle\hat{F}(z_{s}^{t+\nicefrac{{1}}{{2}}})-F(z_{s}^{t+\nicefrac{{1}}{{2}}}),z^{*}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$
$\quad\quad\quad\quad\quad+2\gamma\langle F(z_{s}^{t+\nicefrac{{1}}{{2}}})-F(z^{*}),z^{*}-z_{s}^{t+\nicefrac{{1}}{{2}}}\rangle$
[…]
$\quad\quad\quad\quad\quad-2\gamma\langle F(z_{s}^{t+\nicefrac{{1}}{{2}}})-F(z^{*}),z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\rangle$
[…]
Substituting all the obtained estimates into (14), we arrive at
$0$ $\leqslant\alpha\|z_{s}^{t}-z^{*}\|^{2}-\|z_{s}^{t+1}-z^{*}\|^{2}+(1-\alpha)\|\omega_{s}^{t}-z^{*}\|^{2}-\alpha\|z_{s}^{t+1}-z_{s}^{t}\|^{2}$
$\quad-(1-\alpha)\|z_{s}^{t+1}-\omega_{s}^{t}\|^{2}+\alpha\|z_{s}^{t+1}-z_{s}^{t}\|^{2}-\|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2}$
$\quad+(1-\alpha)\|z_{s}^{t+1}-\omega_{s}^{t}\|^{2}-\alpha\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}\|^{2}-(1-\alpha)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t}\|^{2}$
$\quad+\frac{\gamma^{2}L^{2}}{\tau}\|z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t}\|^{2}+\tau\|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2}+\frac{4\gamma^{2}L^{2}}{\tau}\|z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t}\|^{2}$
$\quad+\tau\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\|^{2}-2\gamma\mu\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\|^{2}.$
By grouping the coefficients of the same terms, we get:
$\|z_{s}^{t+1}-z^{*}\|^{2}$ $\leqslant\alpha\|z_{s}^{t}-z^{*}\|^{2}+(1-\alpha)\|\omega_{s}^{t}-z^{*}\|^{2}$
$\quad+\left(\frac{5\gamma^{2}L^{2}}{\tau}-(1-\alpha)\right)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t}\|^{2}-(1-\tau)\|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2}$
$\quad-(2\gamma\mu-\tau)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\|^{2}-\alpha\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}\|^{2}.$ (15)
Now, we want to estimate the $(2\gamma\mu-\tau)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\|^{2}$ term. To do this we split it into two equal parts. To the first part we add and subtract $\omega_{s}^{t}$, and to the second – $z_{s}^{t}$. After that we use (CS) for both terms:
$-\left(2\gamma\mu-\tau\right)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\|^{2}$ $=-\left(\gamma\mu-\frac{\tau}{2}\right)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\|^{2}-\left(\gamma\mu-\frac{\tau}{2}\right)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z^{*}\|^{2}$
$=-\left(\gamma\mu-\frac{\tau}{2}\right)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t}+\omega_{s}^{t}-z^{*}\|^{2}$
$\quad-\left(\gamma\mu-\frac{\tau}{2}\right)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}+z_{s}^{t}-z^{*}\|^{2}$
$\leqslant\left(\gamma\mu-\frac{\tau}{2}\right)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t}\|^{2}-\left(\frac{\gamma\mu}{2}-\frac{\tau}{4}\right)\|\omega_{s}^{t}-z^{*}\|^{2}$
$\quad+\left(\gamma\mu-\frac{\tau}{2}\right)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}\|^{2}-\left(\frac{\gamma\mu}{2}-\frac{\tau}{4}\right)\|z_{s}^{t}-z^{*}\|^{2}.$
Substituting this into (15),
$\|z_{s}^{t+1}-z^{*}\|^{2}$ $\leqslant\left(\alpha-\frac{\gamma\mu}{2}+\frac{\tau}{4}\right)\|z_{s}^{t}-z^{*}\|^{2}+\left(1-\alpha-\frac{\gamma\mu}{2}+\frac{\tau}{4}\right)\|\omega_{s}^{t}-z^{*}\|^{2}$
$\quad+\left(\frac{5\gamma^{2}L^{2}}{\tau}+\gamma\mu-\frac{\tau}{2}-(1-\alpha)\right)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t}\|^{2}$
$\quad-(1-\tau)\|z_{s}^{t+1}-z_{s}^{t+\nicefrac{{1}}{{2}}}\|^{2}+\left(\gamma\mu-\frac{\tau}{2}-\alpha\right)\|z_{s}^{t+\nicefrac{{1}}{{2}}}-z_{s}^{t}\|^{2}.$
We want to choose parameters such that coefficients before the last three terms would be non-positive. Let us start with $\|z_{s}^{t+\nicefrac{{1}}{{2}}}-\omega_{s}^{t}\|^{2}$ term.
$\text{We pick\penalty 10000\ }\tau$ $=\gamma\mu;$
$\text{We want\penalty 10000\ }1-\alpha$ $\geqslant\frac{5\gamma L^{2}}{\mu}+\frac{\gamma\mu}{2};$
$\text{It is enough for us that\penalty 10000\ }\gamma$ $\leqslant\frac{(1-\alpha)\mu}{6L^{2}}.$
Obviously, with this choice of $\gamma$ and $\alpha$, the last two terms are less than zero. In that way, we obtain:
$\|z_{s}^{t+1}-z^{*}\|^{2}\leqslant\left(\alpha-\frac{\gamma\mu}{4}\right)\|z_{s}^{t}-z^{*}\|^{2}+\left(1-\alpha-\frac{\gamma\mu}{4}\right)\|\omega_{s}^{t}-z^{*}\|^{2}.$
According to the condition for updating the point $\omega_{s}^{t}$,
$\mathbb{E}\|\omega_{s}^{t+1}-z^{*}\|^{2}=p\|z_{s}^{t}-z^{*}\|^{2}+(1-p)\|\omega_{s}^{t}-z^{*}\|^{2}.$
In that way:
$\mathbb{E}\|z_{s}^{t+1}-z^{*}\|^{2}+\frac{1-\alpha}{p}\mathbb{E}\|\omega_{s}^{t+1}-z^{*}\|^{2}$ $\leqslant\left(1-\frac{\gamma\mu}{4}\right)\mathbb{E}\|z_{s}^{t}-z^{*}\|^{2}$
$\quad+\left((1-\alpha)\left(1+\frac{1}{p}-1\right)-\frac{\gamma\mu}{4}\right)\mathbb{E}\|\omega_{s}^{t+1}-z^{*}\|^{2}.$
Now we put $\alpha=1-p$ and obtain:
$\mathbb{E}\|z_{s}^{t+1}-z^{*}\|^{2}+\mathbb{E}\|\omega_{s}^{t+1}-z^{*}\|^{2}\leqslant\left(1-\frac{\gamma\mu}{4}\right)\left(\mathbb{E}\|z_{s}^{t}-z^{*}\|^{2}+\mathbb{E}\|\omega_{s}^{t}-z^{*}\|^{2}\right).$
Denoting $V_{s}^{t}=\mathbb{E}\|z_{s}^{t}-z^{*}\|^{2}+\mathbb{E}\|\omega_{s}^{t}-z^{*}\|^{2}$ and going into recursion over all epochs and iterations, we get:
$V_{S}^{n}\leqslant\left(1-\frac{\gamma\mu}{4}\right)^{T}V_{0}^{0},$
where $T$ is the total number of iterations. ∎
Corollary 2. Suppose that Assumptions 1, 2 hold. Then Algorithm 3.2 with $\gamma\leqslant\frac{(1-\alpha)\mu}{6L^{2}},p=\frac{1}{n}$ and $V_{s}^{t}=\mathbb{E}\|z_{s}^{t}-z^{*}\|^{2}+\mathbb{E}\|\omega_{s}^{t}-z^{*}\|^{2}$, to reach $\varepsilon$-accuracy, where $\varepsilon\sim V_{S}^{n}$, needs
$\mathcal{O}\left(n\frac{L^{2}}{\mu^{2}}\log\left(\frac{1}{\varepsilon}\right)\right)\penalty 10000\ \penalty 10000\ \text{iterations and oracle calls.}$
Proof.
Substituting estimation of $\gamma$ to the result of Theorem 2 we obtain, that method to converge to $\varepsilon$-accuracy, where $\varepsilon=V_{S}^{n}$, needs $\mathcal{O}\left(\frac{L^{2}}{p\mu^{2}}\log\left(\frac{1}{\varepsilon}\right)\right)$ iterations. At the same time each iteration costs $pn+2$ calls to $F_{\pi}$. Thus, we obtain $\mathcal{O}\left(\left(n\frac{L^{2}}{\mu^{2}}+\frac{L^{2}}{p\mu^{2}}\right)\log\left(\frac{1}{\varepsilon}\right)\right)$ oracle complexity. Finally, the optimal choice $p=\frac{1}{n}$ gives $\mathcal{O}\left(n\frac{L^{2}}{\mu^{2}}\log\left(\frac{1}{\varepsilon}\right)\right)$ iteration and oracle complexity. This ends the proof. ∎